# CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time

### v1

@@ Abstract
Pre-trained vision-language models (VLMs), exemplified by CLIP, demonstrate remarkable adaptability across zero-shot classification tasks without additional training. However, their performance diminishes in the presence of domain shifts. In this study, we introduce CLIP Adaptation duRing Test-Time (CLIPArTT), a fully test-time adaptation (TTA) approach for CLIP, which involves automatic text prompts construction during inference for their use as text supervision. Our method employs a unique, minimally invasive text prompt tuning process, wherein multiple predicted classes are aggregated into a single new text prompt, used as pseudo label to re-classify inputs in a transductive manner. Additionally, we pioneer the standardization of TTA benchmarks (e.g., TENT) in the realm of VLMs. Our findings demonstrate that, without requiring additional transformations nor new trainable modules, CLIPArTT enhances performance dynamically across non-corrupted datasets such as CIFAR-10, corrupted datasets like CIFAR-10-C and CIFAR-10.1, alongside synthetic datasets such as VisDA-C. This research underscores the potential for improving VLMs’ adaptability through novel test-time strategies, offering insights for robust performance across varied datasets and environments. The code can be found at: https://github.com/dosowiechi/CLIPArTT.git
@@ 1 Introduction
In this paper, CLIP is contextualized in the setting of Test-Time Adaptation, a challenging yet practical scenario of domain adaptation. In realistic scenarios, a model needs to adapt to new data on-the-fly to cope with unknown distribution shifts, and without using any class supervision. Although an experimental boilerplate was standardized in recent years, CLIP has not been integrated to it yet. Additionally, we introduce a powerful adaptation technique that achieves state-of-the-art performance without a significant computational overhead. Comprehensive studies are performed on three datasets containing different types of domain shifts on several levels of severity, resulting in a total of 155 evaluation scenarios. Our main contributions can then be summarized as follows:
  • We propose a light-weight Test-Time Adaptation method for CLIP, called CLIPArTT, that adapts this VLM to new domains without a considerable computation overhead.
@@
The rest of the paper is organized in the following sections. Section 2 compiles significant related work to contextualize our method. Section 3 introduces CLIPArTT and explains its functionality. Section 4 details the experimental setting, with results exposed in Section 5. Section 6 concludes with a summary of contributions and potential future work.
@@ 2 Related work
An important branch of TTA methods targets batch normalization layers for adaptation, as they can keep key information from the source domain. As the simplest approach, PTBN Nado et al. (2021) addresses the problem by simply recomputing the batch statistics at test-time. TENT Wang et al. (2021) goes further by using entropy minimization on the model’s predictions as feedback, and optimizes the affine parameters in batch normalization layers after one iteration. Entropy minimization has since been a recurrent alternative to update models to target domains without label supervision Goyal et al. (2022); Zhang et al. (2022); Lin et al. (2023); Lee et al. (2023) by increasing the model’s confidence. These approaches, however, rely on image augmentations or large batches to be effective. More recently, methods such as Text-Prompt Tuning (TPT) Shu et al. (2022) have introduced entropy minimization to adapt CLIP, while focusing on learning an adapter for the text prompts. However, this method heavily relies on input augmentations, making it more costly to train. Our method also fine-tunes normalization layers to have a minimal impact on the model’s knowledge and utilizes the confidence of the predictions to improve the text supervision. Moreover, we do not employ any augmentations on the inputs.
@@
Pseudo-labels have also played an important role in previous TTA research. SHOT Liang et al. (2020) employs them with cross-entropy to regularize mutual information between inputs and predictions. The entire feature encoder is optimized in this case. CoTTA Wang et al. (2022) makes use of pseudo-labels in the context of a student-teacher model, where one model processes the original input and the other processes augmented versions of that input. A consistency loss is used to compare both models’ outputs. PAD Wu et al. (2021) improves the quality of pseudo-labels by also augmenting inputs and using voting on the models’ predictions. Instead of simple class pseudo-labels, CLIPArTT improves the text supervision to broaden the likelihood of accurately predicting the right class.
@@
Some other methods utilize less conventional approaches for TTA. LAME Boudiaf et al. (2022) refines the classifier’s predictions in a transductive manner using the similarity between features through the Laplacian. A family of test-time training (TTT) methods have also gained popularity Sun et al. (2020); Liu et al. (2021); Osowiechi et al. (2023); Gandelsman et al. (2022); Vargas Hakim et al. (2023); Osowiechi et al. (2024), where a sub-branch is trained alongside the main network in an unsupervised way, which is later used to update the model. These methods require training the model from scratch on the source domain. As LAME, our method is related to the concept of Laplacian regularization, however we exploit this concept very differently. Whereas LAME employs it to refine predictions in a traditional transductive inference approach where the model is kept as is, our method exploits in a test-time adaptation loss enforcing the consistency between embeddings of related batch samples. Contrary to TTT methods, we do not train the model from scratch with any additional branch.
@@
Conformal Learning.  CLIPArTT is also related to the field of conformal learning, where intervals of confidence for new predictions are derived from previous experience Shafer and Vovk (2008). In a conformal prediction, a given level of certainty is assigned to a set $\mathcal{C}=\{c_{1},...,c_{K}\}$ with the $K$ most plausible classes that an input can belong to Angelopoulos and Bates (2021). We draw inspiration from this concept and build a conformal set of class predictions that can help adapting CLIP towards an accurate top--1 prediction. Our technique, however, stands out by not relying on image transformations such as in Shu et al. (2022); Lai et al. (2023) to filter out predictions.
@@ 3 Methodology
Figure 1: CLIPArTT pipeline overview: 1) Computing predictions from Image-Text Similarity, 2) generating a new text prompt by filtering the top-$K$ class predictions, 3) employing the new prompt to obtain image-to-image and text-to-text similarity scores, used to build a pseudo-label for cross-entropy.
@@ 3.1 CLIP-based classification
$p_{ik}\,=\,\frac{\exp\big{(}\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{k})/\tau\big{)}}{\sum_{j}\exp\big{(}\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{j})/\tau\big{)}},\quad\cos(\mathbf{z},\mathbf{z}^{\prime})=\frac{{\mathbf{z}^{\top}}{\mathbf{z}^{\prime}}}{\|\mathbf{z}\|_{2}\!\cdot\!\|\mathbf{z}^{\prime}\|_{2}},$ (1)
@@
Figure 2: Evolution of the accuracy after several iterations of adaptation at test-time for TENT.
@@ 3.2 Our CLIPArTT method
The proposed CLIPArTT method, illustrated in Figure 1, addresses the above-mentioned limitations using two key insights. The first insight, inspired by conformal learning, is that the correct class is often among the ones with the highest probabilities, although the model’s most confident prediction may not be always correct. This claim is supported by the results in Table 13 (Original), showing that the correct class is within the top-3 predictions 97.79% of the times for CIFAR10 (versus 88.74% within the top-1), and 80.92% of the times for CIFAR100 (versus 61.68% within the top-1). A way to include information about multiple classes could, therefore, help adapting the model at test time. The second insight is that the similarity between the batch samples could be evaluated based on their visual and/or text embeddings. As we will show below, such similarities could be exploited in a strategy related to graph-Laplacian regularization.
@@
Transductive TTA  Next, we design a test-time adaptation loss that accounts for semantic relationships between batch samples. Let $\mathbf{Z}^{v}\in{}^{B\times D}$ and $\hat{\mathbf{Z}}^{t}\in{}^{B\times D}$ denote the normalized visual and instance-specific text embeddings of the samples within the test batch, respectively. We compute an image-to-image similarity matrix, $\mathbf{S}^{v}=\mathbf{Z}^{v}{(\mathbf{Z}^{v})^{\top}}\in[-1,1]^{B\times B}$, and a text-to-text similarity matrix, $\mathbf{S}^{t}=\hat{\mathbf{Z}}^{t}{(\hat{\mathbf{Z}}^{t})^{\top}}\in[-1,1]^{B\times B}$. The former measures the affinity between each pair of samples within the batch in terms of their visual characteristics (shapes, textures, etc.). The latter captures common (or related) classes in the top-$k$ predictions of two samples, since it is computed using the instance-specific multi-class prompts.
@@
Let $\hat{\mathbf{P}}$ denotes the zero-shot prediction matrix using our instance-specific multi-class text prompts:
$\hat{\mathbf{P}}=\text{softmax}\big{(}\mathbf{Z}^{v}{(\hat{\mathbf{Z}}^{t})^{\top}}/\tau\big{)},\quad\hat{p}_{ij}=\frac{\exp\big{(}\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{j})/\tau\big{)}}{\sum_{k}\exp\big{(}\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{k})/\tau\big{)}}$ (4)
@@
Unlike recent approaches Shu et al. (2022), which adapt the CLIP model by learning a text prompt, we instead update the normalization-layer parameters of the visual encoder, which yields a light-weight, computationally efficient TTA method.
3.3 Link to graph-Laplacian regularization
An important element of our TTA method is that it exploits the similarities between pairs of samples within the batch, in terms of the visual features and text prompts representing the top-$k$ classes. In this section, we draw a connection between the proposed method and graph-Laplacian regularization, well-known in the context semi-supervised learning techniques Chapelle et al. (2006) and label propagation Iscen et al. (2019). In our case, the Laplacian regularizer is computed over a set of $B$ nodes with predictions $\mathbf{Z}\in{}^{B\times D}$ as
@@ 3.3 Link to graph-Laplacian regularization
$=\frac{1}{\tau B}\sum_{i}\sum_{j}q_{ij}\Big{(}-\!{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}\,+\,\tau\underbrace{\log\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}_{\text{LogSumExp (LSE)}}\Big{)}$ (9)
@@
The last equality comes from the hypothesis that, in a well trained model, the maximum similarity occurs between the image and embedding of a samples and its corresponding text embedding. Our TTA loss can then be expressed as
$\mathcal{L}_{\mathrm{TTA}}$ $\approx\frac{1}{\tau B}\Big{(}\sum_{i}\sum_{j}q_{ij}\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}\ +\,\ \sum_{i}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}\Big{)}\,+\,\text{const}$ (12)
$=\frac{1}{\tau B}\mathrm{trace}\big{(}{(\mathbf{Z}^{v})^{\top}}\!(\mathbf{L}_{\mathrm{Q}}+\mathbf{I})\,\mathbf{Z}^{t}\big{)}\ +\,\text{const}$ (13)
where $\mathbf{I}$ is the identity matrix. The modified Laplacian $\mathbf{L}_{\mathrm{Q}}+\mathbf{I}$ enforces nodes with a high connection weight ($q_{ij}$) to have similar embeddings, while also avoiding embeddings to collapse into a single vector.

### v2

# CLIPArTT: Adaptation of CLIP to New Domains at Test Time

@@ Abstract
Pre-trained vision-language models (VLMs), exemplified by CLIP, demonstrate remarkable adaptability across zero-shot classification tasks without additional training. However, their performance diminishes in the presence of domain shifts. In this study, we introduce CLIP Adaptation duRing Test-Time (CLIPArTT), a fully test-time adaptation (TTA) approach for CLIP, which involves automatic text prompts construction during inference for their use as text supervision. Our method employs a unique, minimally invasive text prompt tuning process, wherein multiple predicted classes are aggregated into a single new text prompt, used as pseudo label to re-classify inputs in a transductive manner. Additionally, we pioneer the standardization of TTA benchmarks (e.g., TENT) in the realm of VLMs. Our findings demonstrate that, without requiring additional transformations nor new trainable modules, CLIPArTT enhances performance dynamically across non-corrupted datasets such as CIFAR-100, corrupted datasets like CIFAR-100-C and ImageNet-C, alongside synthetic datasets such as VisDA-C. This research underscores the potential for improving VLMs’ adaptability through novel test-time strategies, offering insights for robust performance across varied datasets and environments. The code can be found at: https://github.com/dosowiechi/CLIPArTT.git
@@ 1 Introduction
In this paper, CLIP is contextualized in the setting of Test-Time Adaptation, a challenging yet practical scenario of domain adaptation. In realistic scenarios, a model needs to adapt to new data on-the-fly to cope with unknown distribution shifts, and without using any class supervision. Although an experimental boilerplate was standardized in recent years, CLIP has not been integrated to it yet. Additionally, we introduce a powerful adaptation technique that achieves state-of-the-art performance without a significant computational overhead. Comprehensive studies are performed on multiple datasets containing different types of domain shifts on several levels of severity, resulting in a total of numerous evaluation scenarios. Our main contributions can be summarized as follows:
  • We propose a Test-Time Adaptation method for CLIP, named CLIPArTT, which adapts the VLM by updating normalization-layer parameters. This is achieved by combining multiple classes into a single new text prompt, which is then used as a pseudo-label.
@@ 2 Related work
A key focus in TTA methods is adapting batch normalization layers, which retain important source domain information. PTBN Nado et al. (2021) adjusts batch statistics at test time, while TENT Wang et al. (2021) refines affine parameters using entropy minimization on predictions. Entropy minimization, used in various methods Goyal et al. (2022); Zhang et al. (2022); Lin et al. (2023); Lee et al. (2023), enhances model confidence without label supervision but often depends on image augmentations or large batches. Text-Prompt Tuning (TPT) Shu et al. (2022) applies entropy minimization to CLIP by learning text prompt adapters, though it is costly due to required augmentations. Test-Time Distribution Normalization (TTDN) Zhou et al. (2023) normalizes test data to match the training distribution but needs access to source data or approximations. Our method fine-tunes normalization layers, leverages prediction confidence for text supervision, and avoids input augmentations.
Recent techniques have also sought to efficiently adapt CLIP in a gradient-free manner. CALIP Guo et al. (2023) adapts the visual and text features bidirectionally through a parametric-free attention module, and later combine them to obtain new logits. Although efficient, this method requires the features of the entire test set to perform a hyperparameter search, which limits its generalization to very large datasets. TDA Karmanov et al. (2024) dynamically builds a positive and a negative cache to adapt features through the Tip-Adapter strategy Zhang et al. (2021). This method however requires to find specific hyperparameters for each dataset, as in the original Tip-Adapter approach.
@@
Pseudo-labels have been central to previous TTA methods. SHOT Liang et al. (2020) uses them with cross-entropy to regularize mutual information, optimizing the entire feature encoder. CoTTA Wang et al. (2022) employs pseudo-labels in a student-teacher model with consistency loss between original and augmented inputs. PAD Wu et al. (2021) enhances pseudo-labels by augmenting inputs and voting on predictions. Instead of simple class pseudo-labels, our CLIPArTT method improves the text supervision to broaden the likelihood of accurately predicting the right class.
@@
Other methods take less conventional approaches to TTA. LAME Boudiaf et al. (2022) refines classifier predictions transductively using feature similarity through the Laplacian. Test-time training (TTT) methods Sun et al. (2020); Liu et al. (2021); Osowiechi et al. (2023); Gandelsman et al. (2022); Vargas Hakim et al. (2023) train a sub-branch alongside the main network in an unsupervised manner to update the model, requiring training from scratch on the source domain. Like LAME, our method relates to Laplacian regularization but applies it differently. While LAME refines predictions without altering the model, our method uses it in a test-time adaptation loss to enforce consistency between embeddings of related batch samples. Unlike TTT methods, we do not require additional branches or training from scratch.
@@
Conformal Learning.  CLIPArTT is also related to the field of conformal learning, where intervals of confidence for new predictions are derived from previous experience Shafer and Vovk (2008). In a conformal prediction, a given level of certainty is assigned to a set $\mathcal{C}=\{c_{1},...,c_{K}\}$ with the $K$ most plausible classes that an input can belong to Angelopoulos and Bates (2021). We draw inspiration from this concept and build a conformal set of class predictions that can help adapting CLIP towards an accurate top-1 prediction. Our technique, however, stands out by not relying on image transformations such as in Shu et al. (2022); Lai et al. (2023) to filter out predictions.
@@ 3 Methodology
Figure 1: CLIPArTT pipeline overview: 1) Computing predictions from Image-Text Similarity, 2) generating a new text prompt by filtering the top-$K$ class predictions, 3) with the new prompts, a pseudo-label Q is obtained by averaging the image-to-image and text-to-text similarity scores, while the prediction $\hat{\textbf{P}}$ is computed as the image-to-text similarity. Cross-entropy is then used as the TTA loss.
Figure 2: Top: Example of similarity matrices ($\mathbf{S}^{v}$, $\mathbf{S}^{t}$) and CLIPArTT softmax probabilities ($\mathbf{Q}$) for a batch of 5 examples and using $K=5$ classes. Bottom: a) When using the identity matrix as pseudo-label for contrastive learning, the correct prediction is ambiguous, as the images are forced to both approaching and moving away from the right class. b) CLIPArTT uses soft pseudo-labels that smoothly guides the prediction towards the correct class by reducing the impact of ambiguities in the prompts.
@@ 3.1 CLIP-based classification
$p_{ik}\,=\,\frac{\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{k})/\tau\big{)}}{\sum_{j}\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{j})/\tau\big{)}},\ \ \cos(\mathbf{z},\mathbf{z}^{\prime})=\frac{{\mathbf{z}^{\top}}{\mathbf{z}^{\prime}}}{\|\mathbf{z}\|_{2}\!\cdot\!\|\mathbf{z}^{\prime}\|_{2}},$ (1)
@@
Refer to caption Refer to caption
Figure 3: Evolution of CLIPArTT’s accuracy during test-time adaptation. Left: For different versions of CIFAR10. Right: Compared to TENT on CIFAR10-C.
@@ 3.2 Our CLIPArTT method
The proposed CLIPArTT method (see Figure 1), addresses the above-mentioned limitations using two key insights. The first insight, inspired by conformal learning, is that the correct class is often among the top most probable ones, although the model’s most confident prediction may not be always correct. This claim is supported by the results in Table 1 (first row: Original), showing that the correct class is within the top-3 predictions 97.79% of the times for CIFAR-10 (versus 88.74% within the top-1), and 80.92% of the times for CIFAR-100 (versus 61.68% within the top-1). A way to include information about multiple classes could, therefore, help adapting the model at test time. The second insight is that the similarity between the batch samples could be evaluated based on their visual and/or text embeddings. As we will show below, such similarities could be exploited in a strategy related to Stochastic Neighbor Embedding (SNE) and graph-Laplacian regularization.
@@
Transductive TTA.  Next, we design a test-time adaptation loss that accounts for semantic relationships between batch samples. Let $\mathbf{Z}^{v}\in{}^{B\times D}$ and $\hat{\mathbf{Z}}^{t}\in{}^{B\times D}$ denote the normalized visual and instance-specific text embeddings of the samples within the test batch, respectively. We compute an image-to-image similarity matrix, $\mathbf{S}^{v}=\mathbf{Z}^{v}{(\mathbf{Z}^{v})^{\top}}\in[-1,1]^{B\times B}$, and a text-to-text similarity matrix, $\mathbf{S}^{t}=\hat{\mathbf{Z}}^{t}{(\hat{\mathbf{Z}}^{t})^{\top}}\in[-1,1]^{B\times B}$. The former measures the affinity between each pair of samples within the batch in terms of their visual characteristics (shapes, textures, etc.). The latter captures common (or related) classes in the top-$k$ predictions of two samples, since it is computed using the instance-specific multi-class prompts. As illustrated in Figure 2 (top), a broad range of similarity values are obtained with this approach.
@@
Let $\hat{\mathbf{P}}$ denote the zero-shot prediction matrix using our instance-specific multi-class text prompts:
$\hat{\mathbf{P}}=\text{softmax}\big{(}\mathbf{Z}^{v}{(\hat{\mathbf{Z}}^{t})^{\top}}/\tau\big{)},\,\ \hat{p}_{ij}=\frac{\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{j})/\tau\big{)}}{\sum_{k}\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{k})/\tau\big{)}}$ (4)
@@
Unlike recent approaches like TPT Shu et al. (2022), which adapt CLIP by learning a text prompt, we instead update the normalization-layer parameters of the visual encoder, which yields a light-weight, computationally efficient TTA method. We note that TPT requires to generate multiple augmentations during the forward pass, incurring a substantial overhead.
The mechanism of CLIPArTT is illustrated in Fig. 2. Our soft pseudo-label reduces the risk of increasing the ambiguity of the predictions. Using the identity matrix (i.e., hard pseudo-labels) leads to uncertain class assignments, as the ambiguity of the text prompts would both attract and repel the right class if it is present in the same text prompt for two different images.
3.3 Link to existing techniques
An important element of our TTA method is that it exploits the similarities between pairs of samples within the batch, in terms of the visual features and text prompts representing the top-$k$ classes. In this section, we draw a connection between the proposed method and two well-known techniques in machine learning: Stochastic Neighbor Embedding Hinton and Roweis (2002) and graph-Laplacian regularization Chapelle et al. (2006); Iscen et al. (2019).
Stochastic Neighbor Embedding (SNE).  This popular technique for dimensionality reduction estimates the local probability of a point $\mathbf{x}_{j}$ given a point $\mathbf{x}_{i}$ based on their Euclidean distance $d_{ij}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}$:
$p_{ij}\,=\,\frac{\exp\big{(}-\!d^{2}_{ij}\big{)}}{\sum_{k\neq i}\exp\big{(}-\!d^{2}_{ik}\big{)}}$ (6)
The goal is to find, for each $\mathbf{x}_{i}$, a low-dimensional embedding $\mathbf{y}_{i}$ such that the local probabilities $q_{ij}$ computed using Eq. (6) on these embeddings, is similar to $p_{ij}$. This is achieved by minimizing the row-wise KL divergence between distribution matrices $\mathbf{P}$ and $\mathbf{Q}$. Our CLIPArTT method can be linked to SNE since $\cos(\mathbf{x}_{i},\mathbf{x}_{j})=-\!d^{2}_{ij}+2$ when $\mathbf{x}_{i}$, $\mathbf{x}_{j}$ are unit normalized, and KL divergence between $\mathbf{P}$ and $\mathbf{Q}$ is equal to the cross-entropy between these matrices minus the entropy of $\mathbf{P}$. In summary, our TTA loss in Eq. (5) ensures that the inter-modality (text-to-image) similarities of batch samples are aligned with their intra-modality ones (text-to-text and image-to-image).
Graph-Laplacian regularization  is commonly-used in the context semi-supervised learning techniques Chapelle et al. (2006) and label propagation Iscen et al. (2019). In our case, the Laplacian regularizer is computed over a set of $B$ nodes with predictions $\mathbf{Z}\in{}^{B\times D}$ as:
@@ 3.3 Link to existing techniques
$\,=\,\frac{1}{\tau B}\sum_{i}\sum_{j}q_{ij}\Big{(}-\!{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}$
$\qquad\qquad\quad\ +\,\tau\underbrace{\log\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}_{\text{LogSumExp (LSE)}}\Big{)}$ (10)
@@
The last equality comes from the hypothesis that, in a well trained model, the maximum similarity occurs between the image embedding of a sample and its corresponding text embedding. Our TTA loss can then be expressed as
$\mathcal{L}_{\mathrm{TTA}}$ $\,\approx\,\frac{1}{\tau B}\Big{(}\sum_{i,j}q_{ij}\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}\,+\,\sum_{i}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}\Big{)}\,+\text{\emph{const}}$
$\,=\,\frac{1}{\tau B}\mathrm{trace}\big{(}{(\mathbf{Z}^{v})^{\top}}\!(\mathbf{L}_{\mathrm{Q}}+\mathbf{I})\,\mathbf{Z}^{t}\big{)}\ +\text{\emph{const}}$ (13)
where $\mathbf{I}$ is the identity matrix. The modified Laplacian $\mathbf{L}_{\mathrm{Q}}+\mathbf{I}$ enforces nodes with a high connection weight ($q_{ij}$) to have similar embeddings, while also avoiding embeddings to collapse into a single vector. ∎